Can you help me in translating this statement into equation? "A man has three sums of money invested, one at 4%, one at 5% and one at 3%. His total yearly income from the three investments is P1400. The last investment yields 100 more per year than the other two combined. If he could receive 1% more on each investment, his annual income would be increased by P400. How much does he invested at each rate?
Answer by stanbon(75887)

Can you help me in translating this statement into equation?
"A man has three sums of money invested, one at 4%, one at 5% and one at 3%.
His total yearly income from the three investments is P1400.
---
Interest Equation: 0.04x+0.05y+0.03z = 1400
---
The last investment yields 100 more per year than the other two combined.
Interest Eq: 0.03z = (0.04x+0.05y)+100
-----
If he could receive 1% more on each investment, his annual income would be increased by P400.
Interest Eq: 0.05x + 0.06y + 0.04z = 1400+400
--------------------------------------
How much does he have invested at each rate?
=======================================================
Cheers,
Stan H.